# Active Label Correction for Building LLM-based Modular AI Systems

### v2

@@ Abstract
Large Language Models (LLMs) have been used to build modular AI systems such as HuggingGPT, Microsoft Bing Chat, and more. To improve such systems after deployment using the data collected from human interactions, each module can be replaced by a fine-tuned model but the annotations received from LLMs are low quality. We propose that active label correction can be used to improve the data quality by only examining a fraction of the dataset. In this paper, we analyze the noise in datasets annotated by ChatGPT and study denoising it with human feedback. Our results show that active label correction can lead to oracle performance with feedback on fewer examples than the number of noisy examples in the dataset across three different NLP tasks.
Active Label Correction for Building LLM-based Modular AI Systems
Karan Taneja Georgia Institute of Technology karan.taneja@cc.gatech.edu                        Ashok Goel Georgia Institute of Technology ashok.goel@cc.gatech.edu
@@ 1 Introduction
Large language models (LLMs) such as OpenAI’s ChatGPT OpenAI (2022), Google’s Bard based on LaMDA Thoppilan et al. (2022), and Meta’s LLaMA Touvron et al. (2023) are powerful zero or few-shot learners as they can generalize to a wide range of tasks simply through task explanations or demonstrations without any model fine-tuning Qin et al. (2023). This has led to the development of many modular AI agents that leverage these LLMs through API services to decompose complex tasks and automate workflows. Some popular examples include HuggingGPT Shen et al. (2023), AutoGPT Significant Gravitas (2023), and Bing Chat. Further, many enterprises are building their own conversational AIs based on LLMs as zero-shot learners as well as response-generation models Tellez (2023). These agents rely on several discriminative and generative sub-tasks to achieve the goal specified by the user. For example, one of HuggingGPT’s sub-tasks is to perform model selection from HuggingFace Model Hub based on model descriptions. While producing the final natural language output is a generative task, many sub-tasks like model selection are discriminative.
@@
AI agents based on LLMs work reasonably well in practice, but their robustness is still in question as they are limited by the quality of zero-shot learning of the LLMs that they employ Shen et al. (2023). An error in one module can trigger a cascade effect, propagating the error throughout the modular system. Further, LLMs acting as zero-shot learners cannot be easily fine-tuned for the end tasks in every AI agent for two reasons: (i) a lack of training data, and (ii) the computational cost required to train and maintain an online service for multiple fine-tuned models can be expensive and also has a large carbon footprint. AI developers also cannot simply rely on future improvements in the quality of underlying LLMs as the performance on certain end-tasks can also worsen Chen et al. (2023) likely due to fine-tuning that focuses on the safety and conversationality of LLMs.
In this paper, we discuss a pipeline based on active label correction Rebbapragada et al. (2012) for improving modular AI systems that rely on LLMs as zero-shot learners and we evaluate its efficacy for three NLP tasks. We propose utilizing the data traces containing user inputs and LLM outputs from the initial deployment to learn a smaller substitute model for each module. We employ active label correction which methodically uses human expertise by seeking annotations for examples that are most likely misannotated by LLM. This process encourages leveraging the powerful zero-shot learning capabilities of the LLMs to rapidly deploy modular AI agents and improve their quality as they accumulate data from the real world. We focus on discriminative tasks in our experiments for this paper but the proposed method also extends to generative tasks in principle (see Appendix A).
Theoretically, the optimization problem at our hand is to select the optimal subset of examples that, when corrected by human annotators, will jointly minimize the test error and the cost of correction. Formally, let $\theta$ be the parameters of a machine learning (ML) model $M$ trained using dataset $\mathcal{D}$. $E(\theta)$ be the error of the model over the test set, and $C(\mathcal{D}_{m})$ be the cost of annotating $\mathcal{D}_{m}$ (for example, $C(\mathcal{D}_{m})=|\mathcal{D}_{m}|$). From a dataset $\mathcal{D}$ with noisy annotations, we wish to find the subset $\mathcal{D}_{m}\subseteq\mathcal{D}$ of examples for which we obtain human annotations or corrections to form $\mathcal{D}_{(m)}$. The optimization problem can be written as
$\min_{\mathcal{D}_{m}\subseteq\mathcal{D},\theta}\quad$ $E(\theta)+\eta\cdot C(\mathcal{D}_{m})$
s.t. $\theta=M(\mathcal{D}\setminus\mathcal{D}_{m}\cup\mathcal{D}_{(m)})$ (1)
where $\eta$ is a trade-off factor. This is similar to the machine teaching problem Zhu (2015) for which a closed-form solution does not exist in the general case, and especially in LLMs where model parameters are uninterpretable and the input domain of text is extremely complex. We use active label correction as an iterative greedy approach by alternately (i) seeking corrections for most likely misannotated examples based on predictions from a trained model, and (ii) training the model to learn from the updated annotations.
@@
Our paper makes three main contributions: (i) we propose a pipeline based on active label correction to continually improve modular AI agents while minimizing annotation costs, (ii) we compare performance-cost curves of different ALC methods to study their efficacy on datasets annotated using ChatGPT as a zero-shot learner, and (iii) we present experiments for three NLP tasks that are common to modular AI agents including natural language inference, named entity recognition, and intent classification.
@@ 2 Related Work
LLM-based Modular AI Agents: ChatGPT OpenAI (2022) and LaMDA Thoppilan et al. (2022) are conversational LLMs trained using large amounts of text and fine-tuned with human feedback. ChatGPT has been shown to perform well on many diverse NLP tasks using prompts Qin et al. (2023) though it doesn’t perform at par with fine-tuned models. LLaMa Touvron et al. (2023) is an open-source LLM trained for zero-shot learning and surpasses GPT-3 on many benchmarks. HuggingGPT Shen et al. (2023) is a modular AI agent that utilizes ChatGPT to solve complex tasks like manipulating objects in image, video, text, and audio domains using a large collection of models from HuggingFace [^1] https://huggingface.co/. AutoGPT Significant Gravitas (2023) is another example of an AI agent based on ChatGPT and GPT-4 to solve tasks by planning, goal setting, task prioritization, and task execution. These agents depend on numerous calls to LLMs behind API services and are expensive to run. We propose a pipeline to use data traces from the deployment of AI agents and limited human supervision to improve their performance and to make them cost-effective by using smaller models. We use ChatGPT as our zero-shot learner as it is easily available as an API service and it performs well at following instructions. We found BARD based on LaMDA Thoppilan et al. (2022) to be challenging to use as it consistently ignores the instructions related to the output formatting.
Figure 1: Proposed process for improving LLM-based modular AI agents. A noisy annotated dataset obtained from a zero-shot or few-shot learner-based module is fed to a misannotation predictor to get flagged examples. The flagged examples are corrected by a human annotator and used to update the dataset. This process is performed iteratively until a stopping condition is met and the final dataset is used to train a substitute model.
@@
Semi-supervised Learning and Active Learning: In semi-supervised learning, training data consists of a set of labeled examples and a typically much larger set of unlabeled examples Hady and Schwenker (2013). Weakly supervised learning more generally assumes that annotations are partial, noisy, or provided at a higher level of abstraction Zhou (2018). Self-training is a common method used to train a teacher model on the supervised subset to predict soft pseudo-labels for the unsupervised subset. Then, the supervised subset and pseudo-labeled subset are used to train a student model Karamanolakis et al. (2021). This can be done iteratively and end-to-end to improve pseudo-labels and model parameters Yi and Wu (2019); Tanaka et al. (2018); Yao et al. (2019). On the other hand, active learning is a strategy to select the most informative examples, typically by constructing them or choosing from a pool of unlabeled examples, and querying an oracle for supervision Settles (2009). In our work, we begin with a fully annotated but noisy dataset and evaluate the performance gain and cost reduction associated with iteratively seeking human expertise for improving data and model quality.
@@
Active Label Correction (ALC): ALC, introduced by Rebbapragada et al. (2012), uses the self-training process to clean a noisy training set by iteratively predicting most likely mislabeled examples using a trained model and receiving human feedback to correct them. They find that simple disagreement measures based on the output probabilities work well for identifying mislabeled examples. Robust ALC Kremer et al. (2018) assumes that noise is conditioned on class labels and uses a latent model for the true label. Dual ALC or DALC Li et al. (2022) additionally uses model predictions with high confidence for automatic correction in their work. These works have primarily focused on image classification problems in vision and experimented with synthetic noise. We explore the use of ALC and DALC for correcting noise in ChatGPT’s text annotations. Our proposed pipeline is aimed at leveraging noisy data to improve LLM-based modular AI systems.
@@ 3 Methods
Modular AI agents similar to HuggingGPT Shen et al. (2023) rely on the zero-shot learning power of LLMs to process the input, perform sub-tasks, and generate output. These agents can accumulate large amounts of data including inputs, intermediate outputs of each module, and final outputs through their interactions with users. To evaluate a module that relies on LLMs as a zero-shot learner, one can analyze the inputs and outputs amassed by the module. Figure 1 outlines our pipeline. We propose collecting input-output pairs for each module to create a noisy annotated dataset, and improving data quality using human feedback with the goal of training a replacement model for the zero-shot learner. The dataset formed by input-output pairs is noisy because the output predictions from zero-shot learners have lower quality than that of a human-annotated dataset. Using ALC Rebbapragada et al. (2012), we iteratively predict and flag the examples that are most likely to be noisy or misannotated, and seek correction from human annotators on the flagged examples. This iterative greedy approach to solving Equation 1 is motivated by (i) the goal of minimizing noisy training examples which is a natural proxy for test set error, and (ii) using minimum human feedback by effectively predicting misannotated examples.
@@ 3.1 Misannotation Prediction
The task of misannotation prediction (MP) involves predicting the examples in a dataset that are most likely to contain noise based on the entire noisy dataset. We wish to train a model that learns patterns from the noisy dataset and flags examples that do not conform with these patterns.
@@
Given a noisy dataset $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{n}$ where $(x_{i},y_{i})$ are pairs of inputs and outputs, we train a model $p_{\theta}$ to predict $y^{*}=\arg\max_{y}p_{\theta}(y|x)$. We then obtain predictions on the training dataset itself and store the probabilities $p_{\theta}(y_{i}|x_{i})$ for each example. Similar to Rebbapragada et al. (2012), we suppose that an example is misannotated with probability $m(x_{i},y_{i})=1-p_{\theta}(y_{i}|x_{i})$. In other words, if the probability of example $(x_{i},y_{i})$ is low under the trained model, it is very likely to be misannotated ($m\approx 1$) and vice versa. We refer to this method as ALC.
@@
Intuitively, when the output for a training example is inconsistent with the annotation in the dataset, we can argue that the remaining examples suggest evidence against the given annotation and, therefore, it is likely to be misannotated. Since each example has a minimal impact on the model for a large enough sample size, $p_{\theta}(y_{i}|x_{i})$ approximates as the probability of the model trained with one example removed from the dataset. We study this impact of data size in the experiments and results section.
We also experiment with Random Label Correction or RLC as a baseline where the examples are effectively assigned a random $m$-value i.e. $m(x_{i},y_{i})=\text{Random}(0,1)$.
While ALC employs human supervision for correction, the model predictions can also be a source of correct annotations for a large proportion of examples. Dual Active Label Correction (DALC), introduced by Li et al. (2022), leverages these predictions by using both the model predictions and human feedback to update the dataset. We adapt their method and compare it to ALC by using model predictions to change labels before determining misannotated examples. We accept output label $y^{*}$ as true annotation for an example if $p_{\theta}(y^{*}|x)>\delta$ where $\delta$ is a constant threshold. We refer to this automatic update of annotations based on model predictions as auto-flipping as it requires no human input. For the remaining examples, we use ALC to find the most likely misannotated examples. We shall refer to this method as DALC.
@@
Flagged Examples: After computing $m(x_{i},y_{i})$ $\forall$ $(x_{i},y_{i})\in\mathcal{D}$, we sort $(x_{i},y_{i})\in\mathcal{D}$ from highest to lowest value of $m(x_{i},y_{i})$. We flag $M\in(0,1)$ fraction of examples in $\mathcal{D}$ with the highest $m(.)$ values to be misannotated. The fraction $M$ can be changed to control the number of flagged examples. A low value of $M$ can lead to a low recall as many misannotated examples will be missed in the flagged set while a high $M$ can mean lower precision because many correctly annotated examples may be included in the flagged set. In DALC, we do not flag examples that are auto-flipped.

### v3

# Can Active Label Correction Improve LLM-based Modular AI Systems?

@@ Abstract
Modular AI systems can be developed using LLM-prompts-based modules to minimize deployment time even for complex tasks. However, these systems do not always perform well and improving them using the data traces collected from a deployment remains an open challenge. The data traces contain LLM inputs and outputs, but the annotations from LLMs are noisy. We hypothesize that Active Label Correction (ALC) can be use on the collected data to train smaller task-specific improved models that can replace LLM-based modules. In this paper, we study the noise in three GPT-3.5-annotated datasets and their denoising with human feedback. We also propose a novel method ALC3 that iteratively applies three updates to the training dataset: auto-correction, correction using human feedback and filtering. Our results show that ALC3 can lead to oracle performance with feedback on 17-24% fewer examples than the number of noisy examples in the dataset across three different NLP tasks.
Can Active Label Correction Improve LLM-based Modular AI Systems?
Karan Taneja Georgia Institute of Technology karan.taneja@cc.gatech.edu                        Ashok K. Goel Georgia Institute of Technology ashok.goel@cc.gatech.edu
@@ 1 Introduction
Large language models (LLMs) such as OpenAI’s ChatGPT OpenAI (2022), Google’s Bard based on LaMDA Thoppilan et al. (2022), and Meta’s LLaMA Touvron et al. (2023) are powerful zero or few-shot learners as they can generalize to a wide range of tasks simply through task explanations or demonstrations without any model fine-tuning Qin et al. (2023). This led to the development of many modular AI systems that leverage these LLMs through API services to decompose complex tasks and to automate workflows Shen et al. (2023); Significant Gravitas (2023); Tellez (2023); Taneja et al. (2024). These AI systems rely on several discriminative and generative sub-tasks performed by LLMs, without any fine-tuning, to achieve the desired behavior. LLM-based modular AI systems work reasonably well in practice, but their robustness is still in question because (i) they are limited by the quality of zero-shot learning of the LLMs that they employ Shen et al. (2023), (ii) an error in one module can trigger a cascade effect, propagating the error throughout the modular system, (iii) AI developers cannot rely on future improvements in the quality of underlying LLMs as the performance on certain end-tasks can also worsen Chen et al. (2023) due to the fine-tuning focused on safety and conversationality of LLMs, and (iv) domain shift between LLM training data and deployed environment. Further, LLMs cannot be easily fine-tuned for all sub-tasks in every AI system because (i) there is a lack of training data, and (ii) the computational cost required to train and maintain an online service for multiple fine-tuned models can be expensive and has a large carbon footprint.
Figure 1: Noisy LLM-annotated datasets are collected from deployment of a modular AI system. Active Label Correction (ALC) is used to predict and correct misannotated examples in order to train a replacement model.
To address these challenges, we propose collecting data traces containing LLM inputs and outputs from deployed AI systems, and training smaller task-specific models that can replace LLM calls (see Figure 1). We propose a method based on Active Label Correction (ALC) Rebbapragada et al. (2012) to use human feedback to improve the quality of data obtained from modular AI systems and evaluate its efficacy for three publicly-available NLP datasets. The data traces contain examples with low-quality LLM annotations, but ALC can utilize human expertise in an efficient manner by seeking annotations for examples that are most likely misannotated. We show that synthetic noise models employed by previous works are not adequate to capture noisy LLM annotations which makes it challenging to predict misannotated examples in ALC. The proposed method ALC3 (i) auto-corrects potentially misannotated examples using model predictions, (ii) filters out the confusing examples in addition to (iii) using human feedback. Perfect misannotation prediction on a dataset with $N\%$ noisy examples will require human annotation on $N\%$ of data to clean it. Our results show that the three-step process of ALC3 can lead to oracle test performance with feedback on 17-24% fewer examples than $N\%$ (perfect predictor) for three NLP tasks with different data sizes (5k-100k) and complexity. The proposed method promotes the idea of leveraging the powerful zero-shot learning capabilities of LLMs for rapidly deploying modular AI systems and improving their quality as they accumulate data from the real world. We focus on discriminative tasks but the proposed method also extends to generative tasks in principle (see §A).
@@
We make three main contributions: (i) we propose an ALC-based pipeline and our method ALC3 to continually improve LLM-based modular AI systems while minimizing annotation costs, (ii) we compare different noise models with LLM-annotation noise and the performance-cost curves of different ALC methods for GPT-3.5-annotated datasets, and (iii) we present experiments for three publicly-available NLP datasets with tasks that are common to modular AI systems including natural language inference, named entity recognition, and intent classification.
@@ 2 Related Work
LLM-based Modular AI Systems: ChatGPT OpenAI (2022), LaMDA Thoppilan et al. (2022) are conversational LLMs trained on large amounts of text and fine-tuned with human feedback, and they perform well on diverse NLP tasks but not at par with fine-tuned models Qin et al. (2023). HuggingGPT Shen et al. (2023), AutoGPT Significant Gravitas (2023), Jill Watson Taneja et al. (2024) are examples of modular AI systems based on GPT-3.5 to solve complex tasks like planning, goal setting, manipulating images or texts. These AI systems depend on LLMs as API services and are expensive to run. We propose a pipeline to use data traces from the deployment of AI systems and limited human supervision to improve their performance and to make them cost-effective by using smaller models.
Figure 2: Proposed process for improving LLM-based modular AI systems using ALC3. The inputs and noisy labels from a zero/few-shot learner-based module are used to obtain a trained model. Model predictions on the noisy training dataset are computed for the next three steps. (i) Auto-correction updates the labels where model prediction contradicts the original label with very high confidence. (ii) Human annotation is used to verify and update a fixed number of confusing examples. (iii) Filtering removes some of remaining examples that are deemed noisy based on model predictions. The process is performed iteratively until a stopping condition. Only human annotations are retained after each iteration, iteration two is shown with columns 6, 7, 8, and 9 for illustration.
@@
Semi-supervised, Weakly-supervised and Active Learning: In semi-supervised learning, training data consists of a set of labeled examples and a typically much larger set of unlabeled examples Hady and Schwenker (2013). Weakly supervised learning more generally assumes that annotations are partial, noisy, or provided at a higher level of abstraction Zhou (2018). Self-training is a common method used to train a teacher model on the supervised subset to predict soft pseudo-labels for the unsupervised subset. Then, the supervised subset and pseudo-labeled subset are used to train a student model Karamanolakis et al. (2021); Tanaka et al. (2018); Yi and Wu (2019). Active learning strategy selects the most informative examples, typically by construction or by choosing from a pool of unlabeled examples, and querying an oracle such as human annotator for supervision Settles (2009). Our problem is different from above but lies at the intersection since we begin with a fully-annotated noisy dataset and need to seek human expertise for improving data and model quality.
@@
Active Label Correction (ALC): ALC Rebbapragada et al. (2012) uses self-training to clean a noisy training set by iteratively predicting most likely mislabeled examples using a trained model and receiving human feedback to correct them. Simple disagreement measures based on the output probabilities work well for identifying mislabeled examples. Robust ALC Kremer et al. (2018) assumes that label-conditioned noise and uses a latent model for the true label. Dual ALC or DALC Li et al. (2022) additionally uses model predictions with high confidence for automatic correction. We note that previous works have primarily focused on image classification problems in vision and experimented with synthetic noise. We study and explore the use of ALC for correcting real noise observed in GPT-3.5-annotated datasets.
Human-AI Collaborative Annotation: CoAnnotating Li et al. (2023) is a data annotation process that uses variations in LLM prompts to estimate model uncertainty for deciding allocation to annotators. In our work, we boost data quality towards the goal of continual model improvement to replace LLM-based modules in AI systems. As a result, we use an iterative process where human feedback helps increase model and misannotation prediction quality i.e. better prediction of examples most likely to be misannotated, along with auto-correction and filtering. CoAnnotating is non-iterative and does not leverage fine-tuned models or previously annotated human examples, but relies on numerous LLMs calls for each example. Machine Teaching Taneja et al. (2022a, b) is another framework for human-AI collaborative annotation where annotators work with an ML model and knowledge base to teach concepts to machines in a selective labeling process. In our proposed process, LLMs can be interpreted as a source of knowledge as they provide initial annotations, but the presence of annotations transforms the problem into one of label correction.
@@ 3 Methods
Modular AI systems rely on the zero/few-shot learning power of LLMs to process the input, perform sub-tasks, and generate output. These AI systems can accumulate data traces that include inputs, intermediate outputs of each module, and final outputs through user interactions. To evaluate a module that relies on LLMs as a zero-shot learner, one can analyze the amassed inputs and outputs. We propose collecting these input-output pairs for each module to create noisy annotated datasets, and improving data quality using human feedback with the goal of training a replacement model. The datasets above are assumed to be noisy because LLM annotations have lower quality than human annotations.
ALC Rebbapragada et al. (2012) iteratively predicts and flag the examples that are most likely to be noisy or misannotated, and seeks correction from human annotators on the flagged examples. The iterative approach aims to jointly (i) minimize noise in training examples which is also a proxy for test error, and (ii) minimize human feedback by effectively predicting misannotated examples. Our proposed method ALC3, outlined in Figure 2, uses two additional steps beyond the basic ALC. We first auto-correct the examples where model prediction contradicts the label with very high confidence, leveraging the trained model for labelling, similar to Dual ALC Li et al. (2022). After human annotation, we filter out examples with low label probabilities to reduce the noise in the training dataset. In each iteration, we retain the human annotations, but reset and obtain a new set of auto-corrected and filtered examples. We describe each step in more details in the following subsections.
@@ 3.1 Misannotation Prediction
The task of misannotation prediction (MP) involves predicting the examples that are most likely to be noisy based on the entire noisy dataset. This ensures efficient human annotation as annotators focus on examples most likely to improve the dataset and the trained model. Given a noisy dataset $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{n}$ where $(x_{i},y_{i})$ are pairs of inputs and outputs, we train a model $p_{\theta}$ to predict $y^{*}=\arg\max_{y}p_{\theta}(y|x)$. We then obtain predictions on the $\mathcal{D}$ itself and compute probabilities $p_{\theta}(y_{i}|x_{i})$ for each example. Similar to Rebbapragada et al. (2012), we suppose that an example is misannotated with probability $m(x_{i},y_{i})=1-p_{\theta}(y_{i}|x_{i})$. In other words, if the probability of $(x_{i},y_{i})$ is low under the trained model, it is likely to be misannotated ($m\approx 1$) and vice versa.
@@
Intuitively, when model prediction for a training example is inconsistent with the annotation, we can argue that the remaining examples suggest evidence against the given annotation and, therefore, it is likely to be misannotated. Since each example has a minimal impact on the model for a large enough sample size, $p_{\theta}(y_{i}|x_{i})$ approximates as the probability of the model trained with one example removed from the dataset. We study this impact of data size in the next section.
@@
After computing $m(x_{i},y_{i})$ for all $(x_{i},y_{i})\in\mathcal{D}$, we sort $(x_{i},y_{i})\in\mathcal{D}$ from highest to lowest value of $m(x_{i},y_{i})$. We flag a fraction $M\in(0,1)$ of examples in $\mathcal{D}$ with the highest $m(.)$ values for human annotation. A low value of fraction $M$ leads to a low recall as many misannotated examples are missed while a high $M$ leads to lower precision because many correctly annotated examples are flagged. We refer to this baseline as ALC. We also evaluate Random Label Correction (RLC) as a baseline where the examples are flagged randomly.
3.2 Auto-correction
While ALC only employs human supervision for correction, the model predictions can also be a source of correct annotations. Dual ALC Li et al. (2022) leverages model predictions along with human feedback to update the dataset. Instead of using label-transition matrix for noise modeling, we adapt their method by simply using model predictions to update labels before determining misannotated examples for direct comparison with other ALC methods. We accept output label $y^{*}$ as true annotation for an example if $p_{\theta}(y^{*}|x)>\delta$ where $\delta\lesssim 1$ is a constant threshold. We refer to this as auto-correction since it requires no human input. We use remaining examples to correct the most likely misannotated examples as before and refer to this baseline method as DALC. Since auto-correction is performed before human annotation, we do not flag any examples that are auto-corrected.
3.3 Filtering
In ALC3, after auto-correction and human annotation, we filter out the examples that are likely to be noisy to improve the quality of the training data. Let $m_{\text{flag}}$ be the number of flagged examples using MP and $m_{\text{corr}}$ be the number of corrected examples in the current iteration. Then, MP precision is given by $p_{\text{MP}}=m_{\text{corr}}/m_{\text{flag}}$. In our experiments, we filter out $m_{\text{filter}}$ examples with lowest values of $p_{\theta}(y|x)$ where $m_{\text{filter}}=3m_{\text{corr}}$, a rough approximation of number of examples that will be corrected in the next few iterations.
As $p_{\text{MP}}$ drops with each iteration, effectiveness of filtering in terms of lowering the noise in training dataset reduces. We estimate the noise in every iteration and perform filtering only if $p_{\text{MP}}$ is strictly better than random selection. Therefore, for $k^{th}$ iteration, we use
$m_{\text{filter}}{{}_{,k}}=\begin{cases}3\cdot m_{\text{corr}}{{}_{,k}}&\text{if }p_{\text{MP}}{{}_{,k}}>\eta_{k}\\ 0&\text{otherwise}\end{cases}$
$\eta_{k}=\eta_{o}-\frac{\sum_{j=1}^{k}m_{\text{corr}}{{}_{,j}}}{|\mathcal{D}|}$
where $\eta_{o}$ is the fraction of noisy examples initially which is estimated by annotating a small random subset that is also used as the test set.